OpenStudy (anonymous):

simplify square root of 24

OpenStudy (anonymous):

\[\sqrt{24}\]

OpenStudy (anonymous):

4*6= 24 2*2=4

OpenStudy (anonymous):

so I go \[2\sqrt{6}\]

OpenStudy (anonymous):

yea thats it

OpenStudy (anonymous):

is it like prime factorization from 6th grade?

OpenStudy (anonymous):

alright, I'll do the homework and see if it still makes sense. Thanks

OpenStudy (anonymous):

yes its like prime factorization, and if you see 2 of any number, then you can pull it out